TO: GENO

Supporter Astronomer: Olga

Observer: Nathan, Jon, Chris, Fuyan

Clear sky. Photometric night.

02:18 Collimating the LBT and the seeing is good 0.8". Error when executing the script. 'PROPID' can not be empty

04:34 seeing become worse 1.5" One of the reason is that the target is low, the atmosphere diffraction makes the star elongate.

05:08 Fuyan take over the telescope. The seeing is 1" or so, J10 r26-28 acquisition, takes about 10~min to acquisition. r29-31 science, b11-13 science, (3*20min). There is wired in b29.

06:30 go to J1343 Meet some collimation issues and pointing looks off. Take one 10 mins exposure and stop. Re-collimate the telescope and check the pointing. go back can not find the guide star.. Stop, move on.

07:20 go to bootes J1427 fields

Blue part of the GUI is frozen, the exposure looks like stopped, and IMCS idle, olga look into the log. It seems that the blue side is still taking the exposure. Abort the third exposure. In the command line, type 'abort', 'blue reset', 'red rest' rerun the script, use 'execMODS -e mods.11.J1427.obs', and it does not work. rerun the script, use 'execMODS mods.11.J1427.obs', now it is fine.

3~hour exposure on this mask *11:00 go to the J1433 fields. 1hour and 10 mins exposure

-- FuyanBian - 25 Feb 2012
Topic revision: r5 - 25 Feb 2012, FuyanBian
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback